# Adaptive Prediction Ensemble:

### v1

@@ V-F Yet Another Routing Function
In this section, we aim to evaluate another type of routing function and compare it with the proposed learning-based one that is trained concurrently with the individual predictors. The prediction selection can also be executed by a routing function based on uncertainty estimation. We choose to use the most widely used method, bootstrapping model output variance, as the uncertainty estimation method in the routing function variant. The variant with a routing function based on bootstrapping uncertainty estimation is named APE (BS) in our experiment. Concretely, we use the variance of three MTR prediction outputs as the epistemic uncertainty estimation of the learning-based prediction model, where the three MTR models are randomly initialized and trained on the same training dataset. If the uncertainty estimation surpasses a threshold, then the predictor will discard MTR predictions and choose the constant velocity prediction as the final output, and vice versa.
From Table I, we can see that the performance of APE (BS) is in between the constant velocity model and MTR. The prediction performance of APE (BS) is an interpolation of the two individual predictors. This indicates that the bootstrapping-based uncertainty estimation is noisy and hence inaccurate in this case. In comparison, a learning-based routing function trained concurrently with individual prediction algorithms is more stable and performs better.
@@ VI Conclusions and Discussions
In this work, we tackled the critical challenge of generalizing motion prediction algorithms for autonomous driving across different datasets. The proposed Adaptive Prediction Ensemble framework, incorporating a deep learning expert, a rule-based expert, and a learned routing function, offers a promising solution to improve zero-shot performance. Our experiments demonstrate that by effectively leveraging the strengths of both deep learning and rule-based models, we can achieve substantial gains in prediction accuracy and robustness, especially in challenging out-of-distribution scenarios and long-horizon predictions. While our approach shows promising results, there are several avenues for future explorations. Investigating more sophisticated rule-based models and incorporating additional expert predictors could further enhance the system’s performance. Additionally, exploring different uncertainty estimation techniques for the routing function could lead to more refined decision making.

### v2

@@ V-F Yet Another Routing Function
In this section, we aim to evaluate another type of routing function and compare it with the proposed learning-based one that is trained concurrently with the individual predictors. The prediction selection can also be executed by a routing function based on uncertainty estimation. We choose to use the most widely used method, bootstrapping model output variance, as the uncertainty estimation method in the routing function variant. Concretely, we use the variance of three MTR prediction outputs as the epistemic uncertainty estimation of the learning-based prediction model, where the three MTR models are randomly initialized and trained on the same training dataset. If the uncertainty estimation surpasses a threshold, then the predictor will discard MTR predictions and choose the constant velocity prediction as the final output, and vice versa.
From Table I, we can see that the performance of APE (MTR-bs) is in between the constant velocity model and MTR. The prediction performance of APE (MTR-bs) is an interpolation of the two individual predictors. This indicates that the bootstrapping-based uncertainty estimation is noisy and hence inaccurate in this case. In comparison, a learning-based routing function trained concurrently with individual prediction algorithms is more stable and performs better.
@@ VI Conclusions and Discussions
In this work, we tackled the critical challenge of generalizing motion prediction algorithms for autonomous driving across different datasets. The proposed Adaptive Prediction Ensemble framework, incorporating a deep learning expert, a rule-based expert, and a learned routing function, offers a promising base-model agnostic solution to improve zero-shot performance. Our experiments demonstrate that by effectively leveraging the strengths of both deep learning and rule-based models, we can achieve substantial gains in prediction accuracy and robustness, especially in challenging out-of-distribution scenarios and long-horizon predictions. While our approach shows promising results, there are several avenues for future explorations. Investigating more sophisticated rule-based models and incorporating additional expert predictors could further enhance the system’s performance. Additionally, exploring different uncertainty estimation techniques for the routing function could lead to more refined decision making.